OpenStudy (anonymous):

Divide (8x^4y^3 + 4x^3y^2 - 2x^2y - 12x^2y^4) by -2x^2y.

OpenStudy (igreen):

\(\dfrac{8x^4y^3 + 4x^3y^2 - 2x^2y - 12x^2y^4}{-2x^2y}\) Just regularly divide the -2 into each of the whole numbered terms: \(\dfrac{-4x^4y^3 - 2x^3y^2 + x^2y + 6x^2y^4}{-2x^2y}\)

OpenStudy (igreen):

Now divide the \(x^2\) to every \(x\) term and \(y\) to every \(y\) term. When dividing fractions with like bases, subtract the exponents and keep the base. \(x^2\) has an exponent of 2, and \(y\) has an exponent of 1. So: \(-4x^2y^2 - 2xy + 6y^3\)

OpenStudy (anonymous):

Okay!

OpenStudy (igreen):

Sorry let me correct this: \(\dfrac{8x^4y^3 + 4x^3y^2 - 2x^2y - 12x^2y^4}{-2x^2y}\) Just regularly divide the -2 into each of the whole numbered terms: \(\dfrac{-4x^4y^3 - 2x^3y^2 + x^2y + 6x^2y^4}{x^2y}\) Forgot to remove the \(-2\).

OpenStudy (anonymous):

@iGreen Thanks!!! Now, How would your answer in Part A be affected if the x2 variable in the denominator was just an x? :/